Present value​ (with changing interest​ rates).  Marty has been offered an injury settlement of ​$16 comma 000 payable in 3 years. He wants to know what the present value of the injury settlement is if his opportunity cost is 4.5​%. ​ (The opportunity cost is the interest rate in this​ problem.) What if the opportunity cost is 8​%? What if it is 10.5​%?
Answer #1

Present value=cash flow*Present value of discounting factor(rate%,time period)

At 4.5%:

Present value=16000/1.045^3

=$16000*0.876296604

=$14020.75(Approx).

At 8%:

Present value=16000/1.08^3

=$16000*0.793832241

=$12701.32(Approx)

at 10.5%:

Present value=16000/1.105^3

=$16000*0.741162036

=$11858.59(Approx).